air is being pumped into a spherical balloon at a rate of 10cm^3/min. Determine the rate at which the radius of the balloon is increasing when the radius of the balloon is 20 cm
1
Expert's answer
2019-11-19T13:10:20-0500

"V=(4\/3)\u03c0R^3"

Differentiate w.r.t. t

"\frac{dV}{dt}=(4/3)π3R^2 \frac{dR}{dt}"

"\frac{dV}{dt}=4πR^2 \frac{dR}{dt}"

Given that,

"\frac{dV}{dt}=10cm^3/second"

"R=20cm"

"10=4π(20)^2 \frac{dR}{dt}"

"\frac{dR}{dt}=\frac{10}{4 \pi \cdot (20)^2}=0.0019cm/second"